y=-16x^2+88x+24 how do you solve that with the quadratic equation?
Answer by stanbon(75887)

-16x^2+88x+24 = 0
---
-8(2x^2 - 11x - 3) = 0
a = 2 ; b = -11 ; c = -3
----
Quadratic Formula:
x = [11 +- sqrt(121 - 4*2*-3)]/4
---
x = [11 +- sqrt(145)]/4
---
x = (11/4) + (1/4)sqrt(145) or x = (11/4) - (1/4)sqrt(145)
===============================
Cheers,
Stan H.